### v2

# Nonclassical dynamics of Néel vector and magnetization accompanied by THz and high-harmonic radiation from ultrafast-light-driven NiO antiferromagnet insulator

Abstract
Ultrafast-light-driven strongly correlated antiferromagnetic insulators, such as prototypical NiO with large energy gap eV, have recently attracted experimental attention using photons of both above-gap [K. Gillmeister et al., Nat. Commun. 11, 4095 (2020)] and subgap [H. Qiu et al., Nat. Phys. 17, 388 (2021)] energy. In the latter context, also of great interest to applications, emission of THz radiation is also observed from NiO/Pt bilayers, where heavy metal (HM) Pt introduces strong spin-orbit coupling (SOC). However, in contrast to amply studied spintronic THz emitters using fs laser pulse (fsLP)-driven FM/HM (FM-ferromagnetic metal of conventional type, such as Fe, Ni, Co) bilayers, where ultrafast demagnetization takes place and is directly related to THz emission, microscopic mechanisms of electromagnetic (EM) radiation from NiO/HM bilayer remain obscure as total magnetization of NiO is zero prior to fsLP application. Here we employ the two-orbital Hubbard-Hund-Heisenberg model and study, via numerically exact quantum many-body methods, the dynamics of its Néel vector and nonequilibrium magnetization. This reveals nonclassical (i.e., not describable by Landau-Lifshitz equation) dynamics of Néel vector and nonequilibrium magnetization, changing only in length while not rotating, where the former is substantially reduced in the case above-gap fsLP. Additionally, we compute EM radiation by time-dependence of magnetization, as well as of local charge currents, finding that both contributions are significant in THz frequency range only in NiO with proximity SOC introduced by HM layer. Above THz range, we find integer high-harmonic generation, as well as unusual noninteger harmonics for above-gap fsLP pump.
Introduction.—Pump-probe experiments [1] on strongly correlated antiferromagnetic (AF) insulators (AFI), like prototypical NiO, have revealed [2] intricate fundamental effects, such as coherent long-lived many-body dynamics resulting from a coupling to the AF background where coherent oscillations persists on surprisingly long timescales ps. The fs laser pulse (fsLP) in these experiments has a central frequency that is above-gap between two Hubbard bands [Fig. 2], as it has also been typically the focus of theoretical studies [2, 3]. Theoretical interest also exists to understand quantum tunneling, multiphoton absorption and the so-called “Keldysh crossover” and the ensuing nonlinear doublon-holon pair production in the case of subgap fsLP pump [4, 5, 6, 7]. These findings comprise a panoply of fundamental effects that cannot be found in fsLP-driven conventional band insulators and semiconductors.
The same NiO material driven by subgap fsLP, but in combination with heavy metal (HM) layer like Pt that introduces effects [8] due to strong spin-orbit coupling (SOC), has been recently explored [9] as potential spintronic THz emitter [10, 11, 12, 13, 14, 15]. Isolated ferromagnetic metal (FM) layers (such as Co,Fe,Ni) or FM/HM bilayers have been intensely studied for nearly 30 years in order to understand ultrafast demagnetization [16, 17], as well as THz emission by these systems, which is much stronger in the case of bilayers [10, 11, 12, 13, 14, 18] than in the case of a single FM layer [19, 13, 18]. The fsLP in spintronic experiments typically has a central wavelength of nm, so that its photons have energy centered around eV which is subgap [9, 20] with respect to the separation eV between the two Hubbard bands of NiO. The spintronic phenomena in such experiments on subgap-fsLP-excited NiO have been interpreted [9] by borrowing the standard intuitive picture [10, 11, 12, 13, 14, 18] (for its recent modifications, however, via microscopic theory see Refs. [21, 22, 23]) developed for FM/HM bilayers. That is, an ultrafast spin current is somehow generated that flows from NiO into the HM layer, so that latter can convert it into charge current via the inverse spin Hall effect [24]. The time-dependent charge current is needed to obtain sizable electromagnetic (EM) radiation in the far-field (FF) region [Fig. 1], as well as to interpret [9] the enhancement [13] of the emitted radiation when switching from FM to FM/HM systems. This is because magnetic dipole radiation [19, 13, 18] from time-dependent magnetization is smaller [21] than radiation by a time-dependent charge current [Eqs. (5) vs. (6)]. However, this picture does not explain the microscopic mechanism that generates the assumed spin current in NiO/HM bilayers (only speculation exist thus far [25]), or why the frequency spectrum of such current has features within 0.1–3 THz range that is imprinted in the EM radiation detected experimentally [9]. It is obvious that fsLP will drive electrons into dynamics at its own center frequency , as well as at integer (typically odd [26, 27]) multiples of for sufficiently intense fsLP. Such high-harmonic generation (HHG) in current and EM radiation has been vigorously explored in recent years in diverse quantum materials [28, 29], including strongly correlated ones [6, 7, 30, 31, 26]. Finally, it remains unclear what kind of dynamics is obeyed by the Néel vector (as the difference of two sublattices magnetizations) and magnetization [as the sum of sublattice magnetizations, which is necessarily a nonequilibrium quantity because ], when compared to standard demagnetization [16, 17, 13] of FM layers, where magnetization vector shrinks. In the case of thin FM layers, a rapid and straightforward analysis of the direction of and its magnitude is achieved via [13] magneto-optical Kerr or Faraday effects. In contrast, they do not apply to AFs, so novel ideas have been explored [32] to detect the presumed [33] rotation of the Néel vector.
Thus, developing a microscopic understanding of the response of AFI to subgap fsLP (that is, by starting from a suitable quantum many-body Hamiltonian and using tools of nonequilibrium quantum statistical mechanics), would also help to resolve a number of outstanding issues in AF optospintronics [39]. Note that, specifically for NiO which is a strongly correlated material sharing features of both Mott and charge-transfer insulators [2, 40], X-ray techniques applied as a probe after subgap fsLP pump have revealed [20] possible substantial [1, 3] changes of its electronic structure, such as emergence of midgap states and Hubbard gap reduction persisting on time scales ps. Such phenomena originating from charge dynamics must be taken into account together with local spin dynamics, as they can lead to inextricable complex spin-charge dynamics [2, 41]. In the case of weakly correlated FMs, proper description of spin-charge dynamics is achieved via time-dependent density functional theory (TDDFT) which has provided [42, 43, 44, 45, 46] a most detailed insight into a sequence of fast changing events [47] and their effect [21, 22] on THz emission. However, application [48] of TDDFT to NiO is impeded by intricacies in including a time-dependent [49, 50] Hubbard to properly capture a strong on-site Coulomb interaction that is also dynamical (as opposed to static in conventional equilibrium DFT+U [51] calculations).
In this Letter, we aim to capture all the essential physics of strongly correlated electrons within NiO, in the presence of fsLP and SOC, by employing a two-orbital Hubbard-Hund-Heisenberg (2HHH) model [2] and by simulating its time evolution via numerically exact quantum-many body methods. Its ground AF state is formed by local spins [55] at each site, that are composed of two elemental spins located on two orbitals at that site [Fig. 1], with Hund interaction included between them in order to achieve fixed and stable per atom [56, 57]. In addition, to take into account adjacent nonmagnetic HM layer in THz spintronics experiments [9], we also include SOC of the Rashba type [8, 34] into the 2HHH model of Ref. [2]. Thus, the introduced SOC models proximity effects [8, 36, 37, 38] around the NiO/HM interface, that modifies the electronic structure on the NiO side due to the HM layer (NiO, in turn, modifies bands of HM layer, but we do not include HM layer explicitly due to high computational expense). This model is placed onto a ladder geometry [Fig. 1], allowing for its spectral function [52, 53, 54] to be computed via numerically (quasi)exact simulations using time-dependent density matrix renormalization group (tDMRG) [58, 59, 60, 61] applicable to quasi-one-dimensional lattices. We complement the study with an additional set of numerical simulations using massively parallel exact diagonalization (ED) methods [62] for Hubbard-type systems, implemented within the H package [63, 64]. This technique allows us to access longer times (required for THz radiation calculations) than those possible via tensor network algorithms (like tDMRG) encountering “entanglement barrier” [65, 66].
By considering a multi-orbital ladder, spin-charge separation is largely suppressed, so a plethora of dissipative mechanisms based on electron-spin interaction become active, such as the AF background interaction with doublons and holons [2, 67] or local dissipation caused by the Hund interaction [68, 69, 70]. Our principal results, revealing highly nonclassical dynamics of Néel vector and nonequilibrium magnetization, and the ensuing radiation at both high harmonics of fsLP and in much lower THz range, are given in Figs. 3 and 4. Prior to discussing these results, we introduce useful concepts and notations.
Model and Methods.—The monolayer of NiO is modeled on a tight-binding ladder [Fig. 1] with two orbitals of Ni per site only, albeit O mediated interactions are also included. The same 2HHH model, but without any SOC, was used in Ref. [2] where above-gap pumping of NiO by fsLP was studied experimentally and theoretically. Our Hamiltonian, , is built on top of the model from Ref. [2] by including possible Rashba SOC [8, 34] to describe the NiO/HM bilayer used in THz spintronic experiments [9]. The local terms of account for the Hubbard and Hund physics
Here () is the creation (annihilation) operator of an electron of spin in orbital located at site ; is the corresponding number operator; , , and are the intra-orbital Coulomb, inter-orbital Coulomb, and inter-orbital (or Hund [56, 57]) exchange interaction, respectively; is the onsite chemical potential; and is a magnetic field added [71] on site to lift the degeneracy between spin- and spin- electrons. The Heisenberg exchange interaction between spins at nearest neighbor (NN) sites is given by
where denotes summation over the NN sites; is the electron spin operator expressed using as one of three () Pauli matrices; and in the isotropic case. The kinetic term in tight-binding (TB) Hamiltonian
where is the hopping parameter. An additional TB term is employed to introduce proximity SOC—switched off in NiO case and switched on in NiO/HM case in Figs. 3 and 4—as given by
where denotes row vector of two creation operators and is a direction-dependent matrix hopping [35] with values for horizontal (vertical) bonds. The term is of the Rashba type [34], assumed to originate from proximity to HM layer employed experimentally [9], as often found in DFT calculations [8] on magnet/HM bilayers. Realistic parameters values for NiO are taken from prior first-principles calculations for strongly correlated electrons [72], such as from DFT+U [51] and/or DFT+dynamical mean field theory [73, 74] studies: ; eV; ; and we set and for symmetry reasons [68]. Half-filling is selected by setting the chemical potential to [2]. The magnetic field at site , , as generated by, e.g., a static impurity, induces [71] Néel “checkerboard” order ( and are two NN sites) in the ground state (GS). Nevertheless, the GS at of our simulation retains nonzero entanglement entropy, as witnessed in many recent experiments even at finite temperature [75, 76], so it is not identical to unentangled Néel ket .
The fsLP is introduced via its vector potential [gray line in Figs. 3(e) and 4(e)] of amplitude , which couples to electrons in the form of a Peierls phase [77, 78] multiplying hoppings in Eqs. (3) and (4) by a factor , so and . Here is the dimensionless parameter [79] quantifying the fsLP intensity; is the lattice constant; and fs determines the width of the Gaussian envelope which is initially centered at fs. The center frequency of the fsLP is either eV in Fig. 3, corresponding to a subgap nm wavelength commonly employed in THz spintronic experiments [9, 10, 11, 12, 13, 14, 18]; or eV in Fig. 4 for above-gap fsLP [gaps are shown in Fig. 2]. The electric field of the fsLP is linearly polarized along the -direction (i.e., diagonal to the ladder) in Fig. 1. The time evolution described by on a ladder is obtained via ED [62] by means of the H package [64, 63], where the GS was found using the Lanczos algorithm and the evolution operator is computed via a Taylor expansion considering up to terms. The time step was chosen as .
To compute the EM radiation generated by the charge dynamics, the expectation value of the bond charge current operator [35], from site to site with being a unit matrix, is plugged [80, 81, 23] into the Jefimenko formula [82] for the electric field
The most general Jefimenko formula [82], as proper solution of the Maxwell equations [83], is reorganized [84] above in order to isolate the FF contributions decaying as . We compute radiation at point P in Fig. 1 which is at a distance away from NiO. Here, emphasizes retardation due to relativistic causality; bond currents are assumed to be homogeneous [80, 81, 23] along the path from site to site of length ; is the number of bonds; and we use shorthand notation . In addition, we also compute the electric field of FF radiation by magnetic dipole, i.e., due to the time-dependence of the nonequilibrium magnetization
where indicates the location of site . Time-dependent magnetization and Néel vector are obtained by summing spin expectation values at each site, and , where a “snake-like” enumeration of ladder sites is assumed. Note that other (the - and -components) are vanishingly small.
Results and Discussion.—The usual first take at interpreting experiments on subgap light-driven AFIs, including NiO [85, 86], invokes [85, 87, 88, 86] a direct coupling of light magnetic field (or an effective one due to inverse Faraday effect [86]) local magnetization of AFIs. This leads to classical dynamics of the Néel vector, which rotates without changing its length, in accord with phenomenological theories [33] of Landau-Lifshitz (LL) type [89]. However, limitations of this approach are often found in experiments [86, 90], which is not surprising as light-charge coupling is much stronger [43], so electrons should be explicitly included. But the picture of classical LL dynamics [33] is appealing because it is difficult to develop intuition on how lectrons of AFI, with gapped energy spectrum [Fig. 2], can absorb subgap light to affect its magnetic ordering. Ie case of AF metals, it is easy to envisage (and calculate [81]) howthat fsLP generates photocurrent of conduction electrons, which are then spin-polarized by the magnetic background and exert spin torque [81] onto local magnetization. Its dynamics follows (for weak-intensity fsLP to avoid demagnetization [16, 17]) classical LL dynamics (valid at sufficiently high temperature and for AFIs with [66]). For AFIs, there is no such shortcut and one has to handle complexities of photoexcited Hubbard model [4, 5, 6, 7, 3], which cannot [91], in general, be reduced to studying just their final outcome on magnetic order.
Indeed, our quantum many-body calculations for subgap fsLP-driven NiO reveal highly nonclassical [71, 92] (i.e., outside any description by LL-type theories [33]) dynamics of the Néel vector and magnetization in Fig. 3. That is, both vectors are changing length along the -axis (orthogonal to the ladder in Fig. 1) while not rotating at all. Nonequilibrium magnetization remains zero in plain NiO [Fig. 3(b)]. However, in SO-proximitized NiO [Fig. 3(f)], which is akin to experimentally observed [93] weak ferromagnetism in photodoped Mott insulator with native SOC. The nonzero emits magnetic dipole [Eq. (6)] THz radiation [orange line in Fig. 3(h)], which can (surprisingly, when compared to FM layers [21]) surpass the contribution from [red curve in Fig. 3(h)]. Importantly, THz emission from either of the two sources is significant only only when proximity SOC is present in NiO/HM bilayer [Fig. 3(h)], in full accord with experiments [9]. Thus, our theory explains these experiments without invoking qualitative speculations [9, 25], while showing that concepts borrowed from FM/HM systems (like interlayer spin current and spin-to-charge conversion within HM [10, 11, 13, 14, 18, 22]) are not necessary for THz emission from AFI. The magnetic dipole radiation exhibits even integer HHG [Fig. 3(g)], while odd ones are expected [26] for radiation from , as dictated by symmetry-imposed selection rules of Floquet group theory [94]. Although SOC breaks inversion symmetry, odd integer HHG is preserved [95, 27]. This offers a scheme—detect even HHG in EM radiation from AFI—which directly corroborates how magnetization dynamics in magnets driven by fsLP can be much faster [96, 45] than observed in low-energy transport experiments [25].
For above-gap fsLP in Fig. 4, we find reduction of Néel vector by up to 15%, signifying suppression [3] of AF order in the GS. This is in sharp contrast to its minuscule change for subgap fsLP in Fig. 3. Similarly to subgap fsLP, nonequilibrium magnetization [Fig. 4(f)] emerges only when proximity SOC in switched on. Unlike the subgap fsLP case, charge current and its THz radiation are not significantly enhanced by including proximity SOC [Fig. 4(d) vs. Fig. 4(h)]. The selection rules for HHG remain the same, allowing only odd integer HHG of [Fig. 4(c),(g)] and even ones of [Fig. 4(g)]. Curiously, we also find noninteger HHG of both and , which are beyond standard Floquet theory and its selection rules [94]. Theoretical [26] and experimental [29] reports of noninteger HHG are scarce, and their understanding is in its infancy. For example, they could arise [26] from correlation-driven population of multiple Floquet states (unlike population of a single state used in Floquet group theory [94]), whose exploration we relegate to future studies.
Acknowledgements.
F.G-G. and B.K.N. were supported by the US National Science Foundation (NSF) through the University of Delaware Materials Research Science and Engineering Center, DMR-2011824. The supercomputing time was provided by DARWIN (Delaware Advanced Research Workforce and Innovation Network), which is supported by NSF Grant No. MRI-1919839. A.E.F. was suppported by the US Department of Energy, Office of Basic Energy Sciences under Grant No. DE-SC0022311.References
- Wang et al. [2018] Y. Wang, M. Claassen, C. D. Pemmaraju, C. Jia, B. Moritz, and T. P. Devereaux, Theoretical understanding of photon spectroscopies in correlated materials in and out of equilibrium, Nat. Rev. Mater. 3, 312 (2018).
- Gillmeister et al. [2020] K. Gillmeister, D. Golež, C.-T. Chiang, N. Bittner, Y. Pavlyukh, J. Berakdar, P. Werner, and W. Widdra, Ultrafast coupled charge and spin dynamics in strongly correlated NiO, Nat. Commun. 11, 4095 (2020).
- Wang et al. [2017] Y. Wang, M. Claassen, B. Moritz, and T. P. Devereaux, Producing coherent excitations in pumped Mott antiferromagnetic insulators, Phys. Rev. B 96, 235142 (2017).
- Oka [2012] T. Oka, Nonlinear doublon production in a Mott insulator: Landau-Dykhne method applied to an integrable model, Phys. Rev. B 86, 075148 (2012).
- Shinjo et al. [2022] K. Shinjo, S. Sota, and T. Tohyama, Glassy dynamics of the one-dimensional Mott insulator excited by a strong terahertz pulse, Phys. Rev. Res. 4, L032019 (2022).
- Murakami et al. [2018] Y. Murakami, M. Eckstein, and P. Werner, High-harmonic generation in Mott insulators, Phys. Rev. Lett. 121, 057405 (2018).
- Murakami et al. [2021] Y. Murakami, S. Takayoshi, A. Koga, and P. Werner, High-harmonic generation in one-dimensional Mott insulators, Phys. Rev. B 103, 035110 (2021).
- Grytsyuk et al. [2016] S. Grytsyuk, A. Belabbes, P. M. Haney, H.-W. Lee, K.-J. Lee, M. D. Stiles, U. Schwingenschlögl, and A. Manchon, -asymmetric spin splitting at the interface between transition metal ferromagnets and heavy metals, Phys. Rev. B 93, 174421 (2016).
- Qiu et al. [2021] H. Qiu, L. Zhou, C. Zhang, J. Wu, Y. Tian, S. Cheng, S. Mi, H. Zhao, Q. Zhang, D. Wu, et al., Ultrafast spin current generated from an antiferromagnet, Nat. Phys. 17, 388 (2021).
- Seifert et al. [2016] T. Seifert, S. Jaiswal, U. Martens, J. Hannegan, L. Braun, P. Maldonado, F. Freimuth, A. Kronenberg, J. Henrizi, I. Radu, et al., Efficient metallic spintronic emitters of ultrabroadband terahertz radiation, Nat. Photonics 10, 483 (2016).
- Wu et al. [2017] Y. Wu, M. Elyasi, X. Qiu, M. Chen, Y. Liu, L. Ke, and H. Yang, High-performance THz emitters based on ferromagnetic/nonmagnetic heterostructures, Adv. Mater. 29, 1603031 (2017).
- Jungfleisch et al. [2018] M. B. Jungfleisch, Q. Zhang, W. Zhang, J. E. Pearson, R. D. Schaller, H. Wen, and A. Hoffmann, Control of terahertz emission by ultrafast spin-charge current conversion at Rashba interfaces, Phys. Rev. Lett. 120, 207207 (2018).
- Rouzegar et al. [2022] R. Rouzegar, L. Brandt, L. c. v. Nádvorník, D. A. Reiss, A. L. Chekhov, O. Gueckstock, C. In, M. Wolf, T. S. Seifert, W. Brouwer, et al., Laser-induced terahertz spin transport in magnetic nanostructures arises from the same force as ultrafast demagnetization, Phys. Rev. B 106, 144427 (2022).
- Seifert et al. [2023] T. S. Seifert, D. Go, H. Hayashi, R. Rouzegar, F. Freimuth, K. Ando, Y. Mokrousov, and T. Kampfrath, Time-domain observation of ballistic orbital-angular-momentum currents with giant relaxation length in tungsten, Nat. Nanotechnol. 18, 1132 (2023).
- Leitenstorfer et al. [2023] A. Leitenstorfer, A. S. Moskalenko, T. Kampfrath, J. Kono, E. Castro-Camus, K. Peng, N. Qureshi, D. Turchinovich, K. Tanaka, A. G. Markelz, et al., The 2023 terahertz science and technology roadmap, J. Phys. D 56, 223001 (2023).
- Beaurepaire et al. [1996] E. Beaurepaire, J.-C. Merle, A. Daunois, and J.-Y. Bigot, Ultrafast spin dynamics in ferromagnetic nickel, Phys. Rev. Lett. 76, 4250 (1996).
- Chen et al. [2025] X. Chen, R. Adam, D. E. Bürgler, F. Wang, Z. Lu, L. Pan, S. Heidtfeld, C. Greb, M. Liu, Q. Liu, J. Wang, C. M. Schneider, and D. Cao, Ultrafast demagnetization in ferromagnetic materials: Origins and progress, Phys. Rep. 1102, 1 (2025).
- Liu et al. [2021] Y. Liu, H. Cheng, Y. Xu, P. Vallobra, S. Eimer, X. Zhang, X. Wu, T. Nie, and W. Zhao, Separation of emission mechanisms in spintronic terahertz emitters, Phys. Rev. B 104, 064419 (2021).
- Beaurepaire et al. [2004] E. Beaurepaire, G. M. Turner, S. M. Harrel, M. C. Beard, J.-Y. Bigot, and C. A. Schmuttenmaer, Coherent terahertz emission from ferromagnetic films excited by femtosecond laser pulses, Appl. Phys. Lett. 84, 3465 (2004).
- Wang et al. [2022] X. Wang, R. Y. Engel, I. Vaskivskyi, D. Turenne, V. Shokeen, A. Yaroslavtsev, O. Grånäs, R. Knut, J. O. Schunck, S. Dziarzhytski, et al., Ultrafast manipulation of the NiO antiferromagnetic order via sub-gap optical excitation, Faraday Discuss. 237, 300 (2022).
- Kefayati and Nikolić [2024] A. Kefayati and B. K. Nikolić, Origins of electromagnetic radiation from spintronic terahertz emitters: A time-dependent density functional theory plus Jefimenko equations approach, Phys. Rev. Lett. 133, 136704 (2024).
- Kefayati et al. [2024] A. Kefayati, Y. Ren, M. B. Jungfleisch, L. Gundlach, J. Q. Xiao, and B. K. Nikolic, Deciphering the origin of spin current in spintronic terahertz emitters and its imprint on their electromagnetic radiation via time-dependent density functional theory, arXiv:2410.07360 (2024).
- Varela-Manjarres et al. [2024] J. Varela-Manjarres, A. Kefayati, M. B. Jungfleisch, J. Q. Xiao, and B. K. Nikolić, Charge and spin current pumping by ultrafast demagnetization dynamics, Phys. Rev. B 110, L060410 (2024).
- Saitoh et al. [2006] E. Saitoh, M. Ueda, H. Miyajima, and G. Tatara, Conversion of spin current into charge current at room temperature: Inverse spin-Hall effect, Appl. Phys. Lett. 88, 182509 (2006).
- Han et al. [2023] J. Han, R. Cheng, L. Liu, H. Ohno, and S. Fukami, Coherent antiferromagnetic spintronics, Nat. Mater. 22, 684 (2023).
- Lange et al. [2024] C. S. Lange, T. Hansen, and L. B. Madsen, Noninteger high-order harmonic generation from extended correlated systems, Phys. Rev. A 109, 063103 (2024).
- Tancogne-Dejean et al. [2022] N. Tancogne-Dejean, F. G. Eich, and A. Rubio, Effect of spin-orbit coupling on the high harmonics from the topological Dirac semimetal Na3Bi, npj Comput. Mater. 8, 145 (2022).
- Ghimire and Reis [2018] S. Ghimire and D. A. Reis, High-harmonic generation from solids, Nat. Phys. 15, 10 (2018).
- Schmid et al. [2021] C. P. Schmid, L. Weigl, P. Grössing, V. Junk, C. Gorini, S. Schlauderer, S. Ito, M. Meierhofer, N. Hofmann, D. Afanasiev, et al., Tunable non-integer high-harmonic generation in a topological insulator, Nature 593, 385 (2021).
- Imai et al. [2020] S. Imai, A. Ono, and S. Ishihara, High harmonic generation in a correlated electron system, Phys. Rev. Lett. 124, 157404 (2020).
- Orthodoxou et al. [2021] C. Orthodoxou, A. Zaïr, and G. H. Booth, High harmonic generation in two-dimensional Mott insulators, npj Quantum Mater. 6, 76 (2021).
- Saidl et al. [2017] V. Saidl, P. Němec, P. Wadley, V. Hills, R. P. Campion, V. Novák, K. W. Edmonds, F. Maccherozzi, S. S. Dhesi, B. L. Gallagher, et al., Optical determination of the Néel vector in a CuMnAs thin-film antiferromagnet, Nat. Photonics 11, 91 (2017).
- Galkina and Ivanov [2021] E. G. Galkina and B. A. Ivanov, Phenomenological description of spin dynamics in antiferromagnets: Short history and modern development, Low Temp. Phys. 47, 765 (2021).
- Manchon et al. [2015] A. Manchon, H. C. Koo, J. Nitta, S. M. Frolov, and R. A. Duine, New perspectives for Rashba spin–orbit coupling, Nat. Mater. 14, 871 (2015).
- Nikolić et al. [2006] B. K. Nikolić, L. P. Zârbo, and S. Souma, Imaging mesoscopic spin Hall flow: Spatial distribution of local spin currents and spin densities in and out of multiterminal spin-orbit coupled semiconductor nanostructures, Phys. Rev. B 73, 075303 (2006).
- Marmolejo-Tejada et al. [2017] J. M. Marmolejo-Tejada, K. Dolui, P. Lazić, P.-H. Chang, S. Smidstrup, D. Stradi, K. Stokbro, and B. K. Nikolić, Proximity band structure and spin textures on both sides of topological-insulator/ferromagnetic-metal interface and their charge transport probes, Nano Lett. 17, 5626 (2017).
- Dolui et al. [2020] K. Dolui, U. Bajpai, and B. K. Nikolić, Effective spin-mixing conductance of topological-insulator/ferromagnet and heavy-metal/ferromagnet spin-orbit-coupled interfaces: A first-principles Floquet-nonequilibrium Green function approach, Phys. Rev. Mater. 4, 121201 (2020).
- Žutić et al. [2019] I. Žutić, A. Matos-Abiague, B. Scharf, H. Dery, and K. Belashchenko, Proximitized materials, Mater. Today 22, 85 (2019).
- Němec et al. [2018] P. Němec, M. Fiebig, T. Kampfrath, and A. V. Kimel, Antiferromagnetic opto-spintronics, Nat. Phys 14, 229 (2018).
- Lechermann et al. [2019] F. Lechermann, W. Körner, D. F. Urban, and C. Elsässer, Interplay of charge-transfer and Mott-Hubbard physics approached by an efficient combination of self-interaction correction and dynamical mean-field theory, Phys. Rev. B 100, 115125 (2019).
- Bittner et al. [2018] N. Bittner, D. Golež, H. U. R. Strand, M. Eckstein, and P. Werner, Coupled charge and spin dynamics in a photoexcited doped Mott insulator, Phys. Rev. B 97, 235125 (2018).
- Krieger et al. [2015] K. Krieger, J. K. Dewhurst, P. Elliott, S. Sharma, and E. K. U. Gross, Laser-induced demagnetization at ultrashort time scales: Predictions of TDDFT, J. Chem. Theory Comput. 11, 4870 (2015).
- Chen and Wang [2019] Z. Chen and L.-W. Wang, Role of initial magnetic disorder: A time-dependent ab initio study of ultrafast demagnetization mechanisms, Sci. Adv. 5, eaau8000 (2019).
- Wu et al. [2024] N. Wu, S. Zhang, D. Chen, Y. Wang, and S. Meng, Three-stage ultrafast demagnetization dynamics in a monolayer ferromagnet, Nat. Commun. 15, 2804 (2024).
- Mrudul and Oppeneer [2024] M. S. Mrudul and P. M. Oppeneer, Ab initio investigation of laser-induced ultrafast demagnetization of fept: Intensity dependence and importance of electron coherence, Phys. Rev. B 109, 144418 (2024).
- Shokeen et al. [2017] V. Shokeen, M. Sanchez Piaia, J.-Y. Bigot, T. Müller, P. Elliott, J. K. Dewhurst, S. Sharma, and E. K. U. Gross, Spin flips versus spin transport in nonthermal electrons excited by ultrashort optical pulses in transition metals, Phys. Rev. Lett. 119, 107203 (2017).
- Tengdin et al. [2018] P. Tengdin, W. You, C. Chen, X. Shi, D. Zusin, Y. Zhang, C. Gentry, A. Blonsky, M. Keller, P. M. Oppeneer, et al., Critical behavior within 20 fs drives the out-of-equilibrium laser-induced magnetic phase transition in nickel, Sci. Adv. 4, eaap9744 (2018).
- Dewhurst et al. [2018] J. K. Dewhurst, P. Elliott, S. Shallcross, E. K. U. Gross, and S. Sharma, Laser-induced intersite spin transfer, Nano Lett. 18, 1842 (2018).
- Tancogne-Dejean et al. [2018] N. Tancogne-Dejean, M. A. Sentef, and A. Rubio, Ultrafast modification of Hubbard in a strongly correlated material: Ab initio high-harmonic generation in NiO, Phys. Rev. Lett. 121, 097402 (2018).
- Orhan and O’Regan [2019] O. K. Orhan and D. D. O’Regan, : a critical assessment of the Hubbard correction to exchange-correlation kernels and potentials, Phys. Rev. B 99, 165120 (2019).
- Anisimov et al. [1991] V. I. Anisimov, J. Zaanen, and O. K. Andersen, Band theory and Mott insulators: Hubbard U instead of Stoner I, Phys. Rev. B 44, 943 (1991).
- Zawadzki and Feiguin [2019] K. Zawadzki and A. E. Feiguin, Time- and momentum-resolved tunneling spectroscopy of pump-driven nonthermal excitations in Mott insulators, Phys. Rev. B 100, 195124 (2019).
- Yang and Feiguin [2019] C. Yang and A. E. Feiguin, Spectral function of Mott-insulating Hubbard ladders: From fractionalized excitations to coherent quasiparticles, Phys. Rev. B 99, 235117 (2019).
- Winkler et al. [2017] G. W. Winkler, M. Ganahl, D. Schuricht, H. G. Evertz, and S. Andergassen, Interaction effects in a microscopic quantum wire model with strong spin–orbit interaction, New J. Phys. 19, 063009 (2017).
- Nag et al. [2020] A. Nag, H. C. Robarts, F. Wenzel, J. Li, H. Elnaggar, R.-P. Wang, A. C. Walters, M. García-Fernández, F. M. F. de Groot, M. W. Haverkort, et al., Many-body physics of single and double spin-flip excitations in NiO, Phys. Rev. Lett. 124, 067202 (2020).
- Hoffmann and Blügel [2020] M. Hoffmann and S. Blügel, Systematic derivation of realistic spin models for beyond-Heisenberg solids, Phys. Rev. B 101, 024418 (2020).
- Stadler et al. [2021] K. M. Stadler, G. Kotliar, S.-S. B. Lee, A. Weichselbaum, and J. von Delft, Differentiating Hund from Mott physics in a three-band Hubbard-Hund model: Temperature dependence of spectral, transport, and thermodynamic properties, Phys. Rev. B 104, 115107 (2021).
- White and Feiguin [2004] S. R. White and A. E. Feiguin, Real-time evolution using the density matrix renormalization group, Phys. Rev. Lett. 93, 076401 (2004).
- Daley et al. [2004] A. J. Daley, C. Kollath, U. Schollwöck, and G. Vidal, Time-dependent density-matrix renormalization-group using adaptive effective Hilbert spaces, J. Stat. Mech. 2004, P04005 (2004).
- Schmitteckert [2004] P. Schmitteckert, Nonequilibrium electron transport using the density matrix renormalization group method, Phys. Rev. B 70, 121302 (2004).
- Feiguin [2011] A. E. Feiguin, The density matrix renormalization group and its time-dependent variants, AIP Conf. Proc. 1419, 5 (2011).
- Innerberger et al. [2020] M. Innerberger, P. Worm, P. Prauhart, and A. Kauch, Electron-light interaction in nonequilibrium: exact diagonalization for time-dependent Hubbard hamiltonians, Eur. Phys. J. Plus 135, 922 (2020).
- Ido et al. [2024] K. Ido, M. Kawamura, Y. Motoyama, K. Yoshimi, Y. Yamaji, S. Todo, N. Kawashima, and T. Misawa, Update of H: Newly added functions and methods in versions 2 and 3, Comput. Phys. Commun. 298, 109093 (2024).
- Kawamura et al. [2017] M. Kawamura, K. Yoshimi, T. Misawa, Y. Yamaji, S. Todo, and N. Kawashima, Quantum lattice model solver H, Comput. Phys. Commun. 217, 180 (2017).
- Lerose et al. [2023] A. Lerose, M. Sonner, and D. A. Abanin, Overcoming the entanglement barrier in quantum many-body dynamics via space-time duality, Phys. Rev. B 107, L060305 (2023).
- Garcia-Gaitan and Nikolić [2024] F. Garcia-Gaitan and B. K. Nikolić, Fate of entanglement in magnetism under lindbladian or non-Markovian dynamics and conditions for their transition to Landau-Lifshitz-Gilbert classical dynamics, Phys. Rev. B 109, L180408 (2024).
- Murakami et al. [2023] Y. Murakami, D. Golež, M. Eckstein, and P. Werner, Photo-induced nonequilibrium states in Mott insulators, arXiv:2310.05201 (2023).
- Rincón et al. [2018] J. Rincón, E. Dagotto, and A. E. Feiguin, Photoinduced Hund excitons in the breakdown of a two-orbital Mott insulator, Phys. Rev. B 97, 235104 (2018).
- Lysne et al. [2020a] M. Lysne, Y. Murakami, and P. Werner, Signatures of bosonic excitations in high-harmonic spectra of Mott insulators, Phys. Rev. B 101, 195139 (2020a).
- Strand et al. [2017] H. U. R. Strand, D. Golež, M. Eckstein, and P. Werner, Hund’s coupling driven photocarrier relaxation in the two-band Mott insulator, Phys. Rev. B 96, 165104 (2017).
- Petrović et al. [2021] M. D. Petrović, P. Mondal, A. E. Feiguin, and B. K. Nikolić, Quantum spin torque driven transmutation of an antiferromagnetic Mott insulator, Phys. Rev. Lett. 126, 197202 (2021).
- Held [2007] K. Held, Electronic structure calculations using dynamical mean field theory, Adv. Phys. 56, 829 (2007).
- Kuneš et al. [2007] J. Kuneš, V. I. Anisimov, A. V. Lukoyanov, and D. Vollhardt, Local correlations and hole doping in NiO: A dynamical mean-field study, Phys. Rev. B 75, 165115 (2007).
- Zhang et al. [2019] L. Zhang, P. Staar, A. Kozhevnikov, Y.-P. Wang, J. Trinastic, T. Schulthess, and H.-P. Cheng, calculations of the complex band and tunneling behavior for the transition metal monoxides MnO, FeO, CoO, and NiO, Phys. Rev. B 100, 035104 (2019).
- Scheie et al. [2021] A. Scheie, P. Laurell, A. M. Samarakoon, B. Lake, S. E. Nagler, G. E. Granroth, S. Okamoto, G. Alvarez, and D. A. Tennant, Witnessing entanglement in quantum magnets using neutron scattering, Phys. Rev. B 103, 224434 (2021).
- Laurell et al. [2024] P. Laurell, A. Scheie, E. Dagotto, and D. A. Tennant, Witnessing entanglement and quantum correlations in condensed matter: A review, arXiv:2405.10899 (2024).
- Panati et al. [2003] G. Panati, H. Spohn, and S. Teufel, Effective dynamics for Bloch electrons: Peierls substitution and beyond, Commun. Math. Phys 242, 547 (2003).
- Li et al. [2020] J. Li, D. Golez, G. Mazza, A. J. Millis, A. Georges, and M. Eckstein, Electromagnetic coupling in tight-binding models for strongly correlated light and matter, Phys. Rev. B 101, 205140 (2020).
- Bajpai et al. [2019] U. Bajpai, B. S. Popescu, P. Plecháč, B. K. Nikolić, L. E. F. Foa Torres, H. Ishizuka, and N. Nagaosa, Spatio-temporal dynamics of shift current quantum pumping by femtosecond light pulse, JPhys. Materials 2, 025004 (2019).
- Ridley et al. [2021] M. Ridley, L. Kantorovich, R. van Leeuwen, and R. Tuovinen, Quantum interference and the time-dependent radiation of nanojunctions, Phys. Rev. B 103, 115439 (2021).
- Suresh and Nikolić [2023] A. Suresh and B. K. Nikolić, Quantum classical approach to spin and charge pumping and the ensuing radiation in terahertz spintronics: Example of the ultrafast light-driven weyl antiferromagnet , Phys. Rev. B 107, 174421 (2023).
- Jefimenko [1966] O. D. Jefimenko, Electricity and Magnetism (Appleton Century-Crofts, New York, 1966).
- Griffiths and Heald [1991] D. J. Griffiths and M. A. Heald, Time-dependent generalizations of the Biot–Savart and Coulomb laws, Am. J. Phys. 59, 111 (1991).
- McDonald [1997] K. T. McDonald, The relation between expressions for time-dependent electromagnetic fields given by Jefimenko and by Panofsky and Phillips, Am. J. Phys. 65, 1074 (1997).
- Kampfrath et al. [2011] T. Kampfrath, A. Sell, G. Klatt, A. Pashkin, S. Mährlein, T. Dekorsy, M. Wolf, M. Fiebig, A. Leitenstorfer, and R. Huber, Coherent terahertz control of antiferromagnetic spin waves, Nat. Photonics 5, 31 (2011).
- Satoh et al. [2010] T. Satoh, S.-J. Cho, R. Iida, T. Shimura, K. Kuroda, H. Ueda, Y. Ueda, B. A. Ivanov, F. Nori, and M. Fiebig, Spin oscillations in antiferromagnetic NiO triggered by circularly polarized light, Phys. Rev. Lett. 105, 077402 (2010).
- Kimel et al. [2009] A. V. Kimel, B. A. Ivanov, R. V. Pisarev, P. A. Usachev, A. Kirilyuk, and T. Rasing, Inertia-driven spin switching in antiferromagnets, Nat. Phys. 5, 727 (2009).
- Blank et al. [2023] T. G. H. Blank, K. A. Grishunin, B. A. Ivanov, E. A. Mashkovich, D. Afanasiev, and A. V. Kimel, Empowering control of antiferromagnets by Thz-induced spin coherence, Phys. Rev. Lett. 131, 096701 (2023).
- Landau and Lifshitz [1935] L. D. Landau and E. M. Lifshitz, On the theory of the dispersion of magnetic permeability in ferromagnetic bodies, Phys. Z. Sowjetunion 8, 153 (1935).
- Formisano et al. [2024] F. Formisano, T. T. Gareev, D. I. Khusyainov, A. E. Fedianin, R. M. Dubrovin, P. P. Syrnikov, D. Afanasiev, R. V. Pisarev, A. M. Kalashnikova, J. H. Mentink, et al., Coherent THz spin dynamics in antiferromagnets beyond the approximation of the Néel vector, APL Mater. 12, 011105 (2024).
- Mentink et al. [2015] J. H. Mentink, K. Balzer, and M. Eckstein, Ultrafast and reversible control of the exchange interaction in Mott insulators, Nat. Commun. 6, 6708 (2015).
- Mitrofanov and Urazhdin [2021] A. Mitrofanov and S. Urazhdin, Nonclassical spin transfer effects in an antiferromagnet, Phys. Rev. Lett. 126, 037203 (2021).
- Afanasiev et al. [2019] D. Afanasiev, A. Gatilova, D. J. Groenendijk, B. A. Ivanov, M. Gibert, S. Gariglio, J. Mentink, J. Li, N. Dasari, M. Eckstein, et al., Ultrafast spin dynamics in photodoped spin-orbit Mott insulator , Phys. Rev. X 9, 021020 (2019).
- Neufeld et al. [2019] O. Neufeld, D. Podolsky, and O. Cohen, Floquet group theory and its application to selection rules in harmonic generation, Nat. Commun. 10, 405 (2019).
- Lysne et al. [2020b] M. Lysne, Y. Murakami, M. Schüler, and P. Werner, High-harmonic generation in spin-orbit coupled systems, Phys. Rev. B 102, 081121 (2020b).
- Kimel and Li [2019] A. V. Kimel and M. Li, Writing magnetic memory with ultrashort light pulses, Nat. Rev. Mater. 4, 189 (2019).

### v3

@@
 These findings comprise a panoply of fundamental effects that cannot be found in fsLP-driven conventional band insulators and semiconductors.
-The same NiO material driven by subgap fsLP, but in combination with heavy metal (HM) layer like Pt that introduces effects [8] due to strong spin-orbit coupling (SOC), has been recently explored [9] as potential spintronic THz emitter [10, 11, 12, 13, 14, 15].
+The same NiO material driven by subgap fsLP, but in combination with heavy metal (HM) layer (like Pt,W,Ta) introducing effects [8] due to strong spin-orbit coupling (SOC), have been very recently explored [9, 10] as spintronic THz emitter [11, 12, 13, 14, 15, 16].
 Isolated ferromagnetic metal (FM) layers (such as Co,Fe,Ni) or FM/HM bilayers have been intensely studied for nearly 30 years in order to understand ultrafast demagnetization [17, 18], as well as THz emission by these systems, which is much stronger in the case of bilayers [11, 12, 13, 14, 15, 19] than in the case of a single FM layer [20, 14, 19].
-The fsLP in spintronic experiments typically has a central wavelength of nm, so that its photons have energy centered around eV which is subgap [9, 20] with respect to the separation eV between the two Hubbard bands of NiO.
-The spintronic phenomena in such experiments on subgap-fsLP-excited NiO have been interpreted [9] by borrowing the standard intuitive picture [10, 11, 12, 13, 14, 18] (for its recent modifications, however, via microscopic theory see Refs.
+The fsLP in spintronic experiments typically has a central wavelength of nm, so that its photons have energy centered around eV which is subgap [9, 10, 21] with respect to the separation eV between the two Hubbard bands of NiO.
+The spintronic phenomena in such experiments on subgap-fsLP-excited NiO have been interpreted [9, 10] by borrowing the standard intuitive picture [11, 12, 13, 14, 15, 19] (for its recent modifications, however, via microscopic theory see Refs.
 [22, 23, 24]) developed for FM/HM bilayers.
@@
 The time-dependent charge current is needed to obtain sizable electromagnetic (EM) radiation in the far-field (FF) region [Fig.
-1], as well as to interpret [9] the enhancement [13] of the emitted radiation when switching from FM to FM/HM systems.
-This is because magnetic dipole radiation [19, 13, 18] from time-dependent magnetization is smaller [21] than radiation by a time-dependent charge current [Eqs.
-(5) vs.
-(6)].
-However, this picture does not explain the microscopic mechanism that generates the assumed spin current in NiO/HM bilayers (only speculation exist thus far [25]), or why the frequency spectrum of such current has features within 0.1–3 THz range that is imprinted in the EM radiation detected experimentally [9].
+1], as well as to interpret [9, 10] the enhancement [14] of the emitted radiation when switching from FM to FM/HM systems.
+This is because magnetic dipole radiation [20, 14, 19] from time-dependent magnetization is smaller [22] than radiation by a time-dependent charge current.
+However, this picture does not explain the microscopic mechanism that generates the assumed spin current in NiO/HM bilayers (only speculation exist thus far [26]), or why the frequency spectrum of such current has features within 0.1–3 THz range that is imprinted in the EM radiation detected experimentally [9, 10].
 It is obvious that fsLP will drive electrons into dynamics at its own center frequency , as well as at integer (typically odd [27, 28]) multiples of for sufficiently intense fsLP.
 Such high-harmonic generation (HHG) in current and EM radiation has been vigorously explored in recent years in diverse quantum materials [29, 30], including strongly correlated ones [6, 7, 31, 32, 27].
-Finally, it remains unclear what kind of dynamics is obeyed by the Néel vector (as the difference of two sublattices magnetizations) and magnetization [as the sum of sublattice magnetizations, which is necessarily a nonequilibrium quantity because ], when compared to standard demagnetization [16, 17, 13] of FM layers, where magnetization vector shrinks.
+Finally, it remains unclear what type of dynamics is obeyed by the Néel vector (as the difference of two sublattices magnetizations) and magnetization [as the sum of sublattice magnetizations, which is necessarily a nonequilibrium quantity because ], when compared to standard demagnetization [17, 18, 14] of FM layers, where magnetization vector shrinks.
 In the case of thin FM layers, a rapid and straightforward analysis of the direction of and its magnitude is achieved via [14] magneto-optical Kerr or Faraday effects.
-In contrast, they do not apply to AFs, so novel ideas have been explored [32] to detect the presumed [33] rotation of the Néel vector.
+In contrast, they do not apply to AFs, so novel ideas have been explored [33] to detect the presumed [34, 35, 10] rotation of the Néel vector.
 Thus, developing a microscopic understanding of the response of AFI to subgap fsLP (that is, by starting from a suitable quantum many-body Hamiltonian and using tools of nonequilibrium quantum statistical mechanics), would also help to resolve a number of outstanding issues in AF optospintronics [42].
@@
 1], with Hund interaction included between them in order to achieve fixed and stable per atom [59, 60].
-In addition, to take into account adjacent nonmagnetic HM layer in THz spintronics experiments [9], we also include SOC of the Rashba type [8, 34] into the 2HHH model of Ref.
+In addition, to take into account adjacent nonmagnetic HM layer in THz spintronics experiments [9, 10], we also include SOC of the Rashba type [8, 36] into the 2HHH model of Ref.
 [2].
-Thus, the introduced SOC models proximity effects [8, 36, 37, 38] around the NiO/HM interface, that modifies the electronic structure on the NiO side due to the HM layer (NiO, in turn, modifies bands of HM layer, but we do not include HM layer explicitly due to high computational expense).
+Thus, the introduced SOC models proximity effects [8, 38, 39, 40, 41] around the NiO/HM interface, that modifies the electronic structure on the NiO side due to the HM layer (NiO, in turn, modifies bands of HM layer, but we do not include HM layer explicitly due to high computational expense).
 This model is placed onto a ladder geometry [Fig.
@@
 Prior to discussing these results, we introduce useful concepts and notations.
-Model and Methods.—The monolayer of NiO is modeled on a tight-binding ladder [Fig.
-1] with two orbitals of Ni per site only, albeit O mediated interactions are also included.
+Model and Methods.—The monolayer 111We use monolayer of NiO and do not include HM layer explicitly due to high computational expense of our numerically exact calculations.
+Nevertheless, this limitation is not a drawback for capturing essential features of experiments as they find [10] the highest intensity of emitted THz radiation for the thinnest NiO layer driven by subgap fsLP.
+of NiO is modeled on a tight-binding ladder [Fig.
+1] with two orbitals per site hosting Ni only, while O atom is not modeled explicitly albeit O-mediated interactions areincluded.
 The same 2HHH model, but without any SOC, was used in Ref.
@@
 Our Hamiltonian, , is built on top of the model from Ref.
-[2] by including possible Rashba SOC [8, 34] to describe the NiO/HM bilayer used in THz spintronic experiments [9].
+[2] by including possible Rashba SOC [8, 36] to describe the NiO/HM bilayer used in THz spintronic experiments [9, 10].
 The local terms of account for the Hubbard and Hund physics
@@
 where denotes row vector of two creation operators and is a direction-dependent matrix hopping [37] with values for horizontal (vertical) bonds.
-The term is of the Rashba type [34], assumed to originate from proximity to HM layer employed experimentally [9], as often found in DFT calculations [8] on magnet/HM bilayers.
+The term is of the Rashba type [36], assumed to originate from proximity to HM layer employed experimentally [9, 10], as found in DFT calculations on FM/HM [8] of AFI/HM [40] bilayers.
 Realistic parameters values for NiO are taken from prior first-principles calculations for strongly correlated electrons [76], such as from DFT+U [54] and/or DFT+dynamical mean field theory [77, 78] studies: ; eV; ; and we set and for symmetry reasons [71].
@@
 The magnetic field at site , , as generated by, e.g., a static impurity, induces [75] Néel “checkerboard” order ( and are two NN sites) in the ground state (GS).
-Nevertheless, the GS at of our simulation retains nonzero entanglement entropy, as witnessed in many recent experiments even at finite temperature [75, 76], so it is not identical to unentangled Néel ket .
+Nevertheless, the GS in our simulations retains nonzero entanglement, as witnessed in many recent experiments even at finite temperature [79, 80], so it is not identical to unentangled Néel ket .
 The fsLP is introduced via its vector potential [gray line in Figs.
@@
 The center frequency of the fsLP is either eV in Fig.
-3, corresponding to a subgap nm wavelength commonly employed in THz spintronic experiments [9, 10, 11, 12, 13, 14, 18]; or eV in Fig.
+3, corresponding to a subgap nm wavelength commonly employed in THz spintronic experiments [9, 11, 12, 13, 14, 15, 19, 10]; or eV in Fig.
 4 for above-gap fsLP [gaps are shown in Fig.
@@
 Note that other (the - and -components) are vanishingly small.
-Results and Discussion.—The usual first take at interpreting experiments on subgap light-driven AFIs, including NiO [85, 86], invokes [85, 87, 88, 86] a direct coupling of light magnetic field (or an effective one due to inverse Faraday effect [86]) local magnetization of AFIs.
-This leads to classical dynamics of the Néel vector, which rotates without changing its length, in accord with phenomenological theories [33] of Landau-Lifshitz (LL) type [89].
+Results and Discussion.—The usual first take at interpreting experiments on subgap light-driven AFIs, including NiO [89, 90, 10], invokes [89, 91, 92, 90] a direct coupling of light magnetic field (or an effective one due to inverse Faraday [90] or inverse Cotton-Mouton [10] effects) to local magnetization of AFIs.
+This leads to classical dynamics of the Néel vector, which rotates without changing its length [10], in accord with phenomenological theories [34, 35] of Landau-Lifshitz (LL) type [93].
 However, limitations of this approach are often found in experiments [90, 94], which is not surprising as light-charge coupling is much stronger [46], so electrons should be explicitly included.
-But the picture of classical LL dynamics [33] is appealing because it is difficult to develop intuition on how lectrons of AFI, with gapped energy spectrum [Fig.
-2], can absorb subgap light to affect its magnetic ordering.
+But the picture of classical LL dynamics [34, 35, 10] is appealing because it is difficult to develop intuition on how lectrons of AFI, with gapped energy spectrum [Fig.
+2], absorb subgap light and then affect its magnetic ordering.
 Ie case of AF metals, it is easy to envisage (and calculate [85]) howthat fsLP generates photocurrent of conduction electrons, which are then spin-polarized by the magnetic background and exert spin torque [85] onto local magnetization.
@@
 For AFIs, there is no such shortcut and one has to handle complexities of photoexcited Hubbard model [4, 5, 6, 7, 3], which cannot [95], in general, be reduced to studying just their final outcome on magnetic order.
-Indeed, our quantum many-body calculations for subgap fsLP-driven NiO reveal highly nonclassical [71, 92] (i.e., outside any description by LL-type theories [33]) dynamics of the Néel vector and magnetization in Fig.
+Indeed, our quantum many-body calculations for subgap fsLP-driven NiO reveal highly nonclassical [75, 96] (i.e., outside any description by LL-type theories [34, 10]) dynamics of the Néel vector and magnetization in Fig.
 3.
@@
 3(h)].
-Importantly, THz emission from either of the two sources is significant only only when proximity SOC is present in NiO/HM bilayer [Fig.
-3(h)], in full accord with experiments [9].
-Thus, our theory explains these experiments without invoking qualitative speculations [9, 25], while showing that concepts borrowed from FM/HM systems (like interlayer spin current and spin-to-charge conversion within HM [10, 11, 13, 14, 18, 22]) are not necessary for THz emission from AFI.
+Importantly, THz emission from both of these two sources is significant only only when proximity SOC is present in NiO/HM bilayer [Fig.
+3(h)], in full accord with experiments [9, 10].
+Thus, our theory explains these experiments without invoking qualitative speculations [9, 26, 10], while showing that concepts borrowed from FM/HM systems (like interlayer spin current and spin-to-charge conversion within HM [11, 12, 14, 15, 19, 23]) are not necessary for THz emission from AFI.
 The magnetic dipole radiation exhibits even integer HHG [Fig.
@@
 - Qiu et al. [2021] H. Qiu, L. Zhou, C. Zhang, J. Wu, Y. Tian, S. Cheng, S. Mi, H. Zhao, Q. Zhang, D. Wu, et al., Ultrafast spin current generated from an antiferromagnet, Nat. Phys. 17, 388 (2021).
+- Rongione et al.
+[2023] E.
+Rongione, O.
+Gueckstock, M.
+Mattern, O.
+Gomonay, H.
+Meer, C.
+Schmitt, R.
+Ramos, T.
+Kikkawa, M.
+Mičica, E.
+Saitoh, et al., Emission of coherent thz magnons in an antiferromagnetic insulator triggered by ultrafast spin–phonon interactions, Nat.
+Commun.
+14, 1818 (2023).
 - Seifert et al.
@@
 - Galkina and Ivanov [2021] E. G. Galkina and B. A. Ivanov, Phenomenological description of spin dynamics in antiferromagnets: Short history and modern development, Low Temp. Phys. 47, 765 (2021).
+- Gomonay and Loktev [2010] H. V. Gomonay and V. M. Loktev, Spin transfer and current-induced switching in antiferromagnets, Phys. Rev. B 81, 144427 (2010).
 - Manchon et al. [2015] A. Manchon, H. C. Koo, J. Nitta, S. M. Frolov, and R. A. Duine, New perspectives for Rashba spin–orbit coupling, Nat. Mater. 14, 871 (2015).
@@
 4, 121201 (2020).
+- Dolui et al.
+[2022] K.
+Dolui, A.
+Suresh, and B.
+K.
+Nikolić, Spin pumping from antiferromagnetic insulator spin-orbit-proximitized by adjacent heavy metal: a first-principles Floquet-nonequilibrium Green function study, J.
+Phys.
+Mater.
+5, 034002 (2022).
 - Žutić et al. [2019] I. Žutić, A. Matos-Abiague, B. Scharf, H. Dery, and K. Belashchenko, Proximitized materials, Mater. Today 22, 85 (2019).
@@
 - Strand et al. [2017] H. U. R. Strand, D. Golež, M. Eckstein, and P. Werner, Hund’s coupling driven photocarrier relaxation in the two-band Mott insulator, Phys. Rev. B 96, 165104 (2017).
+- Note [1] We use monolayer of NiO and do not include HM layer explicitly due to high computational expense of our numerically exact calculations.
+Nevertheless, this limitation is not a drawback for capturing essential features of experiments as they find [10] the highest intensity of emitted THz radiation for the thinnest NiO layer driven by subgap fsLP.
 - Petrović et al.

### v4

@@
 Above THz range, we find integer high-harmonic generation, as well as unusual noninteger harmonics for above-gap fsLP pump.
-Introduction.—Pump-probe experiments [1] on strongly correlated antiferromagnetic (AF) insulators (AFI), like prototypical NiO, have revealed [2] intricate fundamental effects, such as coherent long-lived many-body dynamics resulting from a coupling to the AF background where coherent oscillations persists on surprisingly long timescales ps.
-The fs laser pulse (fsLP) in these experiments has a central frequency that is above-gap between two Hubbard bands [Fig.
-2], as it has also been typically the focus of theoretical studies [2, 3].
+Introduction.—Pump-probe experiments [1] with strongly correlated antiferromagnetic (AF) insulators (AFI), like prototypical NiO [2], reveal [3] exotic effects interweaving nonequilibrium many-body physics and quantum coherence that can persist on surprisingly long timescales (such as ps [2]) due to a large Mott gap [Fig.
+2] providing protection against fast thermalization and heating.
+The fs laser pulse (fsLP) in these experiments and typical theoretical studies [2, 4] has a central frequency that is above-gap between two Hubbard bands [Fig.
+2].
 Theoretical interest also exists to understand quantum tunneling, multiphoton absorption and the so-called “Keldysh crossover” and the ensuing nonlinear doublon-holon pair production in the case of subgap fsLP pump [5, 6, 7, 8].
-These findings comprise a panoply of fundamental effects that cannot be found in fsLP-driven conventional band insulators and semiconductors.
+Such panoply of complex nonequilibrium many-body states [3] cannot be found in fsLP-driven conventional band insulators and semiconductors, where single particle [9] description suffices.
 The same NiO material driven by subgap fsLP, but in combination with heavy metal (HM) layer (like Pt,W,Ta) introducing effects [10] due to strong spin-orbit coupling (SOC), have been very recently explored [11, 12] as spintronic THz emitter [13, 14, 15, 16, 17, 18].
@@
 It is obvious that fsLP will drive electrons into dynamics at its own center frequency , as well as at integer (typically odd [29, 30]) multiples of for sufficiently intense fsLP.
-Such high-harmonic generation (HHG) in current and EM radiation has been vigorously explored in recent years in diverse quantum materials [29, 30], including strongly correlated ones [6, 7, 31, 32, 27].
+Such high-harmonic generation (HHG) in current and EM radiation has been vigorously explored in recent years in diverse quantum materials [31, 32], including strongly correlated ones [7, 8, 33, 34, 29, 3].
 Finally, it remains unclear what type of dynamics is obeyed by the Néel vector (as the difference of two sublattices magnetizations) and magnetization [as the sum of sublattice magnetizations, which is necessarily a nonequilibrium quantity because ], when compared to standard demagnetization [19, 20, 16] of FM layers, where magnetization vector shrinks.
@@
 In the case of weakly correlated FMs, proper description of spin-charge dynamics is achieved via time-dependent density functional theory (TDDFT) which has provided [47, 48, 49, 50, 51] a most detailed insight into a sequence of fast changing events [52] and their effect [24, 25] on THz emission.
-However, application [51] of TDDFT to NiO is impeded by intricacies in including a time-dependent [52, 53] Hubbard to properly capture a strong on-site Coulomb interaction that is also dynamical (as opposed to static in conventional equilibrium DFT+U [54] calculations).
+However, application [53] of TDDFT to NiO is impeded by intricacies in including a time-dependent [54, 55, 3] Hubbard to properly capture a strong on-site Coulomb interaction that is also dynamical (as opposed to static in conventional equilibrium DFT+U [56] calculations).
 In this Letter, we aim to capture all the essential physics of strongly correlated electrons within NiO, in the presence of fsLP and SOC, by employing a two-orbital Hubbard-Hund-Heisenberg (2HHH) model [2] and by simulating its time evolution via numerically exact quantum-many body methods.
@@
 By considering a multi-orbital ladder, spin-charge separation is largely suppressed, so a plethora of dissipative mechanisms based on electron-spin interaction become active, such as the AF background interaction with doublons and holons [2, 3] or local dissipation caused by the Hund interaction [72, 73, 74].
-Our principal results, revealing highly nonclassical dynamics of Néel vector and nonequilibrium magnetization, and the ensuing radiation at both high harmonics of fsLP and in much lower THz range, are given in Figs.
+Our principal results, revealing highly nonclassical dynamics of Néel vector and nonequilibrium magnetization, and the ensuing radiation at both HHG of fsLP and in much lower THz range, are given in Figs.
 3 and 4.
@@
 11, 4095 (2020).
+- Murakami et al. [2023] Y. Murakami, D. Golež, M. Eckstein, and P. Werner, Photo-induced nonequilibrium states in Mott insulators, arXiv:2310.05201 (2023).
 - Wang et al. [2017] Y. Wang, M. Claassen, B. Moritz, and T. P. Devereaux, Producing coherent excitations in pumped Mott antiferromagnetic insulators, Phys. Rev. B 96, 235142 (2017).
@@
 - Murakami et al. [2021] Y. Murakami, S. Takayoshi, A. Koga, and P. Werner, High-harmonic generation in one-dimensional Mott insulators, Phys. Rev. B 103, 035110 (2021).
+- Bajpai et al.
+[2019] U.
+Bajpai, B.
+S.
+Popescu, P.
+Plecháč, B.
+K.
+Nikolić, L.
+E.
+F.
+Foa Torres, H.
+Ishizuka, and N.
+Nagaosa, Spatio-temporal dynamics of shift current quantum pumping by femtosecond light pulse, JPhys.
+Materials 2, 025004 (2019).
 - Grytsyuk et al.
@@
 B 109, L180408 (2024).
-- Murakami et al. [2023] Y. Murakami, D. Golež, M. Eckstein, and P. Werner, Photo-induced nonequilibrium states in Mott insulators, arXiv:2310.05201 (2023).
 - Rincón et al. [2018] J. Rincón, E. Dagotto, and A. E. Feiguin, Photoinduced Hund excitons in the breakdown of a two-orbital Mott insulator, Phys. Rev. B 97, 235104 (2018).
@@
 B 101, 205140 (2020).
-- Bajpai et al.
-[2019] U.
-Bajpai, B.
-S.
-Popescu, P.
-Plecháč, B.
-K.
-Nikolić, L.
-E.
-F.
-Foa Torres, H.
-Ishizuka, and N.
-Nagaosa, Spatio-temporal dynamics of shift current quantum pumping by femtosecond light pulse, JPhys.
-Materials 2, 025004 (2019).
 - Ridley et al. [2021] M. Ridley, L. Kantorovich, R. van Leeuwen, and R. Tuovinen, Quantum interference and the time-dependent radiation of nanojunctions, Phys. Rev. B 103, 115439 (2021).